1. A steel tie rod on bridge be made withstand a pull of 3500lbs. find the

diameter of the rod assuming a factor of safety of 2.5 and ultimate stress of

5700psi.


( Ans d=1.398 in. )



1
Expert's answer
2022-03-31T02:46:01-0400

"A=\frac{KP}{\sigma}"

"=\frac{2.5*3500}{5700}=\frac{175}{114}"

"A=\frac{\pi}{4}*D^2"

"\frac{175}{114}=\frac{\pi}{4}*D^2"

"D^2=\frac{175*4}{114*\pi}=1.9545"

D=1.398in.